Reports 1-1 of 1 Clear search Modify search
AdV-ALS (Green sources)
gosselin, spinicelli, derossi - 18:32 Monday 17 August 2020 (49235) Print this report
INJ Pick-off for ALS

Today we worked on the pick-off from INJ to be sent to the end buidings.

We had about 40 mW at the input of the fiber in the laser lab.
About 35 mW at the output of the fiber in DAQ room.
About 7-8 mW at the input of both telecom fibers towards WEB and NEB.
About 4 mW at the output of the telecom fiber (which is coherent with the 1dB/km attenuation of those fibers at 1064nm)
About 3.5 mW at the level of the rack at NE.

All those measures have been done with a photometer which has a doubtful calibration (to be checked). We did the last measurment with another photometer and got 2.4 mW. The budget loss can anyway be calculated from those measurments.

We added an in-line polarizer at the output of that fiber and get about 1 mW at the output (with the calibrated photometer).
We plugged it to a photodiode connected to the channel ALS_NEB_PD_IR_MONI. We did not change the offset and the gain of the channel but here are the calibration factors: offset 32 mV, gain 0.5 mW / V.
We will monitor this signal for few days in order to better know the possible variation of power. In that way we can better set the gain of the SOA to stay in its working interval ( 2 mW - 20 mW at the output).

To be noticed:
- We had troubles with the fiber connection between CEB and NEB ( CEB- - NEB 24 SM OS2). It seems that at least the ports 13 and 14 are not working properly. We then tested ports 1 and 20 that are working fine. Our guess is that the bundle of fibers from 13 to 18 is not working, to be checked by the company. 
- We also checked with an IR card that the beam is actually reaching WE (port 13 of the patch panel CEB- - WEB 24 SM OS2). We will measure the power later.

Comments to this report:
gosselin, spinicelli - 15:58 Wednesday 26 August 2020 (49267) Print this report

The calibration that we gave was actually wrong by a factor 2. It is not 0.5 mW/V but 0.25 mW/V.
Attached to this entry you can see the trend over the last 9 days. If we except the moment when the SL was unlocking (due to IMC unlocks), the variation over that period of time was of about a factor 3, between 2 V and 6 V, ie 0.6 mW to 1.5 mW. 
The power is actually greater and more stable than what has been measured in 2018 (log entry

This pick-off will be used to seed the 2 W amplifier at NE. Since this amplifier has a security when its seed is too low, we were worrying about the moments when the SL unlocks.
Speaking with the manufacturer this morning, they told us about an "autostart mode", that automatically switches OFF when the seed is too low and switches ON again when it is high enough. 
We are waiting for their instruction to implement that mode and be able to start long term acquisition of the ALS set-up at NE with the pick-off coming from INJ.

Images attached to this comment
gosselin, spinicelli, melo - 16:37 Tuesday 15 September 2020 (49425) Print this report

Today while installing the fibered laser in DAQ room instead of the pick-off from the injection we realized that the input fiber connector of the splitter box was actually broken inside.
This explains why there was more losses than expected between the 35 mW at the output of the fiber in DAQ room and the 7-8 mW at the input of both telecom fibers.
We replaced the faulty connector (and also the others) and measured about 15 mW instead of the 7-8 mW.
When plugging back the fiber from the injection, we will have to tune the gain of the SOA in order to adapt to this power.

Search Help